# On the complexity of matrix Putinar’s Positivstellensatz

### v1

@@
Abstract.
This paper studies the complexity of matrix Putinar’s Positivstellensatz on the semialgebraic set that is given by the polynomial matrix inequality. Under the archimedeanness, we prove a polynomial bound on degrees of terms appearing in the representation of matrix Putinar’s Positivstellensatz. Estimates on the exponent and constant are given. As a byproduct, a polynomial bound on the convergence rate of matrix sum-of-squares relaxations is obtained, which resolves an open question raised by Dinh and Pham. When the constraining set is unbounded (the archimedeanness fails), we also prove a similar bound for the matrix version of Putinar–Vasilescu’s Positivstellensatz by exploiting homogenization techniques.
Key words and phrases:
polynomial optimization, matrix optimization, Moment-SOS relaxations, Positivstellensatz
2020 Mathematics Subject Classification:
@@ 1. Introduction
Then, it reduces to the classical Positivstellensatz in real algebraic geometry and polynomial optimization. A polynomial $p$ is said to be a sum of squares (SOS) if $p=p_{1}^{2}+\dots+p_{t}^{2}$ for polynomials $p_{1},\dots,p_{t}$. Schmüdgen [44] proved that if $K$ is compact and $F>0$ on $K$, then there exist SOS polynomials $\sigma_{\alpha}$ such that
@@
Under the archimedeanness (a condition slightly stronger than the compactness, referring to Subsection 2.2), Putinar [38] showed that if $F>0$ on $K$, then there exist SOS polynomials $\sigma_{0},\sigma_{1},\dots,\sigma_{m}$ such that
(1.4) $F=\sigma_{0}+\sigma_{1}g_{1}+\sigma_{m}g_{m},$
which is referred to as Putinar’s Positivstellensatz in the literature.
The complexities of the above scalar Positivstellensatze have been studied extensively. Suppose that $K\subseteq(-1,1)^{n}$ and the degree of $F$ is $d$. For the degree $d$, denote
@@
For a polynomial $p=\sum\limits_{\alpha\in\mathbb{N}_{d}^{n}}p_{\alpha}\frac{|\alpha|!}{\alpha_{1}!\cdots\alpha_{n}!}x^{\alpha}$, denote the norm
$\|p\|=\max\{p_{\alpha}:\alpha\in\mathbb{N}_{d}^{n}\}.$
It was shown in [46] that if $F\geq F_{\min}$ on $K$, then for
@@
there exist SOS polynomials $\sigma_{\alpha}$ with $\deg(\sigma_{\alpha}g_{1}^{\alpha_{1}}\cdots g_{m}^{\alpha_{m}})\leq k$ such that (1.3) holds. Here, $c$ is a positive constant depending on constraining polynomials $g_{1},\dots,g_{m}$. Based on the above bound, Nie and Schweighofer [35] proved that under the archimedeanness, there exists a constant $c>0$ depending on $g_{1},\dots,g_{m}$ such that if
@@
then (1.4) holds for some SOS polynomials $\sigma_{1}$, $\dots$, $\sigma_{m}$ with $\deg(\sigma_{1}g_{1})\leq k$ , $\dots$, $\deg(\sigma_{m}g_{m})\leq k$. Recently, Baldi and Mourrain improved the exponent bound (1.5) to a polynomial bound. It was shown in [2] that the same result holds if
@@
In the above, $\gamma$ is a positive constant depending on $n$, $g_{1},\dots,g_{m}$, $L$ is the Lojasiewicz exponent associated with $g_{1},\dots,g_{m}$ and $\|F\|_{\max}$ is the max norm on the cube $[-1,1]^{n}$. The bound (1.6) has been further improved in [3] by removing the dependency of the exponent on the dimension $n$ and providing estimates on the constant $\gamma$. There exist better degree bounds when the set $K$ is specified, seeing [1, 4, 12, 48, 49, 50]. We also refer to [26] when $F,g_{1},\dots,g_{m}$ admit the correlative sparse sparsity.
Based on Putinar’s Positivstellensatz, Lasserre [28] introduced the Moment-SOS hierarchy for solving polynomial optimization and proved its asymptotic convergence under the archimedeanness. The above degree bounds give estimates on convergence rates of the Moment-SOS hierarchy directly. We also refer to [8, 19, 20, 21, 32, 33, 41] for the finite convergence of the Moment-SOS hierarchy under some regular assumptions.
When $F$, $G$ are general polynomial matrices, a matrix version of Putinar’s Positivstellensatz was provided in [25, 43]. Under the archimedeanness, it was proved that if $F\succ 0$ on $K$, then there exist polynomial matrices $P_{i}\in\mathbb{R}[x]^{m\times\ell}$ ($i=1,\dots,t$) and an SOS matrix $S$ such that
(1.7) $F=S+P_{1}^{T}GP_{1}+\cdots P_{t}^{T}GP_{t}.$
A polynomial matrix $S(x)$ is SOS if $S(x)=H(x)^{T}H(x)$ for some polynomial matrix $H(x)$. For the special case that $G$ is diagonal (i.e., $K$ is as in (1.2)), the exponential bound (1.5) was generalized to the matrix representation (1.7) in [14]. When $G$ is a general polynomial matrix, the complexity of matrix Putinar’s Positivstellensatz (1.7) is open, to the best of author’s knowledge. We also refer to [7, 24, 45] for other types of SOS representations for polynomial matrices.
@@
This paper studies the complexity of matrix Putinar’s Positivstellensatz. Throughout the paper, we assume that
@@
unless it is specified. Denote the scaled simplex
@@
Let $\|F\|_{B}$ be the Bernstein norm of $F$ on the scaled simplex $\bar{\Delta}$ (referring to Section 3 for details). The set of all polynomial matrices with representations (1.7) is the quadratic module of $\ell$$-\ell$ polynomial matrices generated by $G(x)$ in $\mathbb{R}[x]^{\ell\times\ell}$, for which we denote by $\mbox{QM}[G]^{\ell}$. The set $\mbox{QM}[G]^{\ell}$ is archimedean if there exists a scalar $r>0$ such that $(r-\|x\|_{2}^{2})I_{\ell}\in\mathrm{QM}[G]^{\ell}$.
The following is our main result. It generalizes the recent results of scalar Putinar’s Positivstellensatz as in [2, 3] to the matrix case and gives the first degree bound for matrix Putinar’s Positivstellensatz.
@@
Let $K$ be as in (1.1) with $m\geq 2$. Suppose $F(x)$ is an $\ell$$-\ell$ symmetric polynomial matrix of degree $d$, and $(1-\|x\|_{2}^{2})\cdot I_{\ell}\in\mbox{QM}[G]^{\ell}$. If $F(x)\succeq F_{\min}\cdot I_{m}\succ 0$ on $K$, then for
@@
(1.12) $F=S+P_{1}^{T}GP_{1}+\cdots P_{t}^{T}GP_{t}.$
In the above, $\kappa,\eta$ are respectively Lojasiewicz constant and exponent as in (5.3), and $C>0$ is a constant independent of $F$, $G$.
Positivity certificate (1.12) can be applied to construct matrix SOS relaxations for solving polynomial matrix optimization [15, 43]. As a byproduct of Theorem 1.1, we can get a polynomial bound on the convergence rate of matrix SOS relaxations (see Theorem 5.5), which resolves an open question raised in [10].
Remark. The assumption that $(1-\|x\|_{2}^{2})\cdot I_{\ell}\in\mbox{QM}[G]^{\ell}$ is not serious. Under the archimedeanness of $\mbox{QM}[G]^{\ell}$, i.e., $(r-\|x\|_{2}^{2})I_{\ell}\in\mathrm{QM}[G]^{\ell}$ for some $r>0$, we have that $(1-\|x\|_{2}^{2})I_{\ell}\in\mathrm{QM}[G(\sqrt{r}x)]^{\ell}$. Theorem 1.1 remains true with $F,G$ replaced by $F(\sqrt{r}x)$, $G(\sqrt{r}x)$.
When $K$ is unbounded, the Putinar type SOS representation (1.7) may not generally hold. For the scalar case (i.e., $F$ is a scalar polynomial and $K$ is described by $g_{1}(x)\geq 0,\dots,g_{m}(x)\geq 0$), it was shown in [39, 40] that if the highest degree homogeneous term of $F$ is positive definite in $\mathbb{R}^{n}$ and $F>0$ on $K$, then there exists a degree $k$ such that $(1+\|x\|_{2}^{2})^{k}F$ has a representation (1.4), which is referred to as Putinar-Vasilescu’s Positivstellensatz. A polynomial complexity was proved for this type of representation by Mai and Magron [31]. Recently, a matrix version of Putinar–Vasilescu’s Positivstellensatz was established in [11].
In the following, we prove a similar bound as in Theorem 1.1 for matrix Putinar–Vasilescu’s Positivstellensatz. Let $F^{hom}$ denote the highest degree homogeneous term of $F$. The homogenization of $F$ is denoted by $\widetilde{F}$, i.e., $\widetilde{F}(\tilde{x}):=x_{0}^{\deg(F)}F(x/x_{0})$ for $\tilde{x}:=(x_{0},x)$. The notation $\lambda_{\min}(A)$ denotes the smallest eigenvalue of the matrix $A$.
@@
(1.13) $(1+\|x\|_{2}^{2})^{k}F=S+P_{1}^{T}GP_{1}+\cdots P_{t}^{T}GP_{t}.$
In the above, $\kappa$, $\eta$ are respectively Lojasiewicz constant and exponent as in (5.8), $C>0$ is a constant independent of $F$, $G$ and $\widetilde{F}_{\min}$ is the optimal value of the optimization
@@
Outline of proofs
@@
Step 1. When $F(x)$ is an $\ell$$-\ell$ symmetric polynomial matrix and $K$ is described by finitely many scalar polynomial inequalities (i.e., $K$ is as in (1.2) ), we establish a polynomial bound on degrees of terms appearing in matrix Putinar’s Positivstellensatz (see Theorem 3.3). It improves the exponential bound given by Helton and Nie [14]. The proof is provided in Section 3.
@@
Step 3. Finally, we prove Theorem 1.1 by applying the main result in Step 1 (i.e, Theorem 3.3) to $F$ and equivalent scalar polynomial inequalities for $K$ obtained in Step 2. We refer to 5.
@@
2. Preliminaries
2.1. Notations
The symbol $\mathbb{N}$ (resp., $\mathbb{R}$) denotes the set of nonnegative integers (resp., real numbers). Let $\mathbb{R}[x]:=\mathbb{R}\left[x_{1},\ldots,x_{n}\right]$ be the ring of polynomials in $x:=\left(x_{1},\ldots,x_{n}\right)$ with real coefficients and let $\mathbb{R}[x]_{d}$ be the ring of polynomials with degrees $\leq d$. For $\alpha=(\alpha_{1},\ldots,\alpha_{n})\in\mathbb{N}^{n}$ and a degree $d$, denote $|\alpha|\coloneqq\alpha_{1}+\cdots+\alpha_{n}$,
$\mathbb{N}_{d}^{n}:=\left\{\alpha\in\mathbb{N}^{n}\mid|\alpha|\leq d\right\}.$
For a polynomial $p,\deg(p)$ denotes its total degree. The degree of a polynomial matrix $P=(P_{ij}(x))\in\mathbb{R}[x]^{\ell_{1}\times\ell_{2}}$ is denoted by $\deg(P)$, i.e.,

### v2

# On the complexity of matrix Putinar’s Positivstellensätz

@@
Abstract
This paper studies the complexity of matrix Putinar’s Positivstellensätz on the semialgebraic set that is given by the polynomial matrix inequality. When the quadratic module generated by the constrained polynomial matrix is Archimedean, we prove a polynomial bound on the degrees of terms appearing in the representation of matrix Putinar’s Positivstellensätz. Estimates on the exponent and constant are given. As a byproduct, a polynomial bound on the convergence rate of matrix sum-of-squares relaxations is obtained, which resolves an open question raised by Dinh and Pham. When the constraining set is unbounded, we also prove a similar bound for the matrix version of Putinar–Vasilescu’s Positivstellensätz by exploiting homogenization techniques.
keywords:
Polynomial matrix, polynomial optimization, matrix Putinar’s Positivstellensätz, matrix sum-of-squares relaxations
{MSCcodes}
@@ 1 Introduction
Then, it reduces to the classical Positivstellensätz in real algebraic geometry and polynomial optimization. A polynomial $p$ is said to be a sum of squares (SOS) if $p=p_{1}^{2}+\dots+p_{t}^{2}$ for polynomials $p_{1},\dots,p_{t}$. Schmüdgen [49] proved that if $K$ is compact and $F>0$ on $K$, then there exist SOS polynomials $\sigma_{\alpha}$ such that
@@
When the quadratic module of scalar polynomials $\mbox{QM}[G]$ generated by $G$ is Archimedean (a condition slightly stronger than compactness, refer to Section 2.2), Putinar [42] showed that if $F>0$ on $K$, then there exist SOS polynomials $\sigma_{0},\sigma_{1},\dots,\sigma_{m}$ such that
(1.4) $F=\sigma_{0}+\sigma_{1}g_{1}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+\cdots+}\sigma_{m}g_{m},$
which is referred to as Putinar’s Positivstellensätz in the literature.
The complexities of the above scalar Positivstellensätze have been studied extensively. Suppose that $K\subseteq(-1,1)^{n}$ and the degree of $F$ is $d$. For the degree $d$, denote
@@
For a polynomial $p=\sum\limits_{\alpha\in\mathbb{N}_{d}^{n}}p_{\alpha}\frac{|\alpha|!}{\alpha_{1}!\cdots\alpha_{n}!}x^{\alpha}$, denote the norm $\|p\|=\max\{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}|p_{\alpha}|}:\alpha\in\mathbb{N}_{d}^{n}\}.$ It was shown in [51] that if ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}F\geq F_{\min}>0}$ on $K$, then for
@@
there exist SOS polynomials $\sigma_{\alpha}$ with $\deg(\sigma_{\alpha}g_{1}^{\alpha_{1}}\cdots g_{m}^{\alpha_{m}})\leq k$ such that (1.3) holds. Here, $c$ is a positive constant that depends on the constraining polynomials $g_{1},\dots,g_{m}$. Based on the above bound, Nie and Schweighofer [39] proved that if $\mbox{QM}[G]$ is Archimedean, there exists a constant $c>0$, depending on $g_{1},\dots,g_{m}$, such that if
@@
then (1.4) holds for some SOS polynomials $\sigma_{0}$, $\dots$, $\sigma_{m}$ with $\deg(\sigma_{0})\leq k$, $\deg(\sigma_{i}g_{i})\leq k$ $(i=1,\dots,m)$. Recently, Baldi and Mourrain improved the exponent bound (1.5) to a polynomial bound. It was shown in [2] that the same result holds if
@@
In the above, $\gamma$ is a positive constant depending on $n$, $g_{1},\dots,g_{m}$, $L$ is the Łojasiewicz exponent associated with $g_{1},\dots,g_{m}$ and $\|F\|_{\max}$ is the max norm on the cube $[-1,1]^{n}$. The bound (1.6) has been further improved in [3] by removing the dependency of the exponent on the dimension $n$ and providing estimates on the constant $\gamma$. There exist better degree bounds when the set $K$ is specified, see [1, 4, 12, 53, 54, 55]. We also refer to [29] when $F,g_{1},\dots,g_{m}$ admit the correlative sparse sparsity.
Based on Putinar’s Positivstellensätz, Lasserre [31] introduced the Moment-SOS hierarchy for solving polynomial optimization and proved its asymptotic convergence. The above degree bounds give estimates on the convergence rates of the Moment-SOS hierarchy directly. We also refer to [8, 20, 21, 23, 36, 37, 46] for the finite convergence of the Moment-SOS hierarchy under some regular assumptions.
When $F$, $G$ are general polynomial matrices, a matrix version of Putinar’s Positivstellensätz was provided in [28, 48]. Let $\mathbb{R}[x]:=\mathbb{R}\left[x_{1},\ldots,x_{n}\right]$ be the ring of polynomials in $x:=\left(x_{1},\ldots,x_{n}\right)$ with real coefficients. A polynomial matrix $S(x)$ is SOS if $S(x)=H(x)^{T}H(x)$ for some polynomial matrix $H(x)$. The set of all $\ell$$-\ell$ SOS polynomial matrices in $x$ is denoted as $\Sigma[x]^{\ell}$. The quadratic module of $\ell$$-\ell$ polynomial matrices generated by $G(x)$ is the set
$\mbox{QM}[G]^{\ell}:=\left\{S+\sum_{i=1}^{t}P_{i}^{T}GP_{i}\mid t\in\mathbb{N},~{}S\in\Sigma[x]^{\ell},~{}P_{i}\in\mathbb{R}[x]^{m\times\ell}\right\}.$
The set $\mbox{QM}[G]^{\ell}$ is Archimedean if there exists a scalar $r>0$ such that $(r-\|x\|_{2}^{2})\cdot I_{\ell}\in\mathrm{QM}[G]^{\ell}$. When the quadratic module $\mbox{QM}[G]^{\ell}$ is Archimedean, it was proved in [28, 48] that if $F\succ 0$ on $K$, then we have that $F\in\mbox{QM}[G]^{\ell}$. For the case that $F$ is a polynomial matrix and $G$ is diagonal (i.e., $K$ is as in (1.2)), the exponential bound (1.5) was generalized to the matrix case in [15]. There are specific estimates on degree bounds when $G$ is a simple set defined by scalar polynomial inequalities. We refer to [12] for the sphere case and refer to [55] for the boolean hypercube case. When $G$ is a general polynomial matrix, the complexity of matrix Putinar’s Positivstellensätz is open, to the best of the author’s knowledge. We also refer to [7, 27, 50] for other types of SOS representations for polynomial matrices.
@@
This paper studies the complexity of matrix Putinar’s Positivstellensätz. Throu-
ghout the paper, we assume that
@@
unless otherwise specified. Denote the scaled simplex
@@
Let $\|F\|_{B}$ be the Bernstein norm of $F$ on the scaled simplex $\bar{\Delta}$ (see Section 3 for details).
The following is our main result. It generalizes the recent results of scalar Putinar’s Positivstellensätz, as shown in [2, 3], to the matrix case and provides the first degree bound for matrix Putinar’s Positivstellensätz.
@@
Let $K$ be as in (1.1) with $m\geq 2$. Suppose $F(x)$ is an $\ell$$-\ell$ symmetric polynomial matrix of degree $d$, and $(1-\|x\|_{2}^{2})\cdot I_{\ell}\in\mbox{QM}[G]^{\ell}$. If $F(x)\succeq F_{\min}\cdot{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}I_{\ell}}\succ 0$ on $K$, then for
@@
(1.11) $F=S+P_{1}^{T}GP_{1}+\cdots+P_{t}^{T}GP_{t}.$
In the above, $\kappa,\eta$ are respectively the Łojasiewicz constant and exponent in (5.3), and $C>0$ is a constant independent of $F$, $G$. Furthermore, the Łojasiewicz exponent can be estimated as
$\eta\leq(3^{m-1}d_{G}+1)(3^{m}d_{G})^{n+\theta(m)-2}.$
The positivity certificate (1.11) can be applied to construct matrix SOS relaxations for solving polynomial matrix optimization [16, 48]. As a byproduct of Theorem 1.1, we can obtain a polynomial bound on the convergence rate of matrix SOS relaxations (see Theorem 5.7), which resolves an open question raised in [10].
When $F(x)$ is a scalar polynomial and $G(x)$ is diagonal with diagonal entries $g_{1}(x),\dots,g_{m}(x)\in\mathbb{R}[x]$, it was shown in [3] that if $F\geq F_{\min}>0$ on $K$, then $f$ admits a representation as in (1.4) for
$k\geq C\cdot 2^{\eta}\cdot n^{2}md_{G}^{6}d^{2\eta}\kappa^{7}\left(\frac{\|F\|_{B}}{F_{\min}}\right)^{7\eta+3},$
where $\kappa,\eta$ are respectively the Łojasiewicz constant and exponent, and $C>0$ is a constant independent of $F$, $G$. Comparing this with the bound as in Theorem 1.1, one important additional factor is $3^{6(m-1)}\theta(m)^{3}$. This comes from the estimates on the quantity and degree bounds of scalar polynomials that can be used to describe the polynomial matrix inequality $G(x)\succeq 0$. Please see Section 4 for details. The other additional factors such as ${8}^{7\eta}$, $d^{14\eta}$ are artifacts of the proof technique. They primarily arise from the degree estimates of $h_{i}g_{i}$ in Theorem 3.5. Since we want the constant $C$ to be independent of $F$ and $G$, we retain all these factors in Theorem 1.1.
Remark. The assumption that $(1-\|x\|_{2}^{2})\cdot I_{\ell}\in\mbox{QM}[G]^{\ell}$ is not serious. When the quadratic module $\mbox{QM}[G]^{\ell}$ is Archimedean, i.e., $(r-\|x\|_{2}^{2})\cdot I_{\ell}\in\mathrm{QM}[G]^{\ell}$ for some $r>0$, we have that $(1-\|x\|_{2}^{2})\cdot I_{\ell}\in\mathrm{QM}[G(\sqrt{r}x)]^{\ell}$. Theorem 1.1 remains true with $F,G$ replaced by $F(\sqrt{r}x)$, $G(\sqrt{r}x)$.
When $K$ is unbounded, the Putinar-type SOS representation may not hold. For the scalar case (i.e., $F\in\mathbb{R}[x]$ and $K$ is as in (1.2)), it was shown in [43, 44] that if the highest degree homogeneous term of $F$ is positive definite in $\mathbb{R}^{n}$ and $F>0$ on $K$, then there exists a degree $k$ such that $(1+\|x\|_{2}^{2})^{k}F$ has a representation (1.4), which is referred to as Putinar-Vasilescu’s Positivstellensätz. A polynomial complexity for this type of representation was proved by Mai and Magron [35]. Recently, a matrix version of Putinar–Vasilescu’s Positivstellensätz was established in [11].
In the following, we state a bound similar to that in Theorem 1.1 for matrix Putinar–Vasilescu’s Positivstellensätz. Let $F^{hom}$ denote the highest degree homogeneous term of $F$. The homogenization of $F$ is denoted by $\widetilde{F}$, i.e., $\widetilde{F}(\tilde{x}):=x_{0}^{\deg(F)}F(x/x_{0})$ for $\tilde{x}:=(x_{0},x)$. The notation $\lambda_{\min}(A)$ denotes the smallest eigenvalue of the matrix $A$.
@@
(1.12) $(1+\|x\|_{2}^{2})^{k}F=S+P_{1}^{T}GP_{1}+\cdots+P_{t}^{T}GP_{t}.$
In the above, $\kappa$, $\eta$ are respectively the Łojasiewicz constant and exponent in (5.8), $C>0$ is a constant independent of $F$, $G$, and $\widetilde{F}_{\min}$ is the optimal value of the optimization
@@
Furthermore, the Łojasiewicz exponent can be estimated as
$\eta\leq(2\cdot 3^{m-1}\lceil d_{G}/2\rceil+1)(2\cdot 3^{m}\lceil d_{G}/2\rceil)^{n+\theta(m)-1}.$
For unbounded optimization, the positivity of $F^{\hom}$ is also crucial for obtaining SOS-type representations because it reflects the behavior of $F$ at infinity (e.g., see [22]). The quantity $\widetilde{F}_{\min}$ is the optimal value of the homogenized optimization (1.15). It is an important quantity for evaluating the positivity of $F$ on the unbounded set $K$, as it reflects not only the behavior of $F$ on $K$ but also its behavior at infinity. Suppose that $F(x)\succeq F_{\min}\cdot I_{\ell}\succ 0$ on $K$, $F^{hom}\succeq F_{\min}^{\prime}\cdot I_{\ell}\succ 0$ on $\mathbb{R}^{n}$, and $\tilde{x}=(x_{0},x)\in\mathbb{R}^{n+1}$ is a feasible point of (1.15). One can see that $x/x_{0}\in K$ and $\widetilde{F}(\tilde{x})=x_{0}^{d}F(x/x_{0})\succeq x_{0}^{d}F_{\min}\cdot I_{\ell}$ for all feasible $\tilde{x}$ with $x_{0}\neq 0$, and $\widetilde{F}(\tilde{x})\succeq F^{\prime}_{\min}\cdot I_{\ell}$ for all feasible $\tilde{x}$ with $x_{0}=0$. However, it is impossible to bound $\widetilde{F}_{\min}$ from below with $F^{hom}$ and $F_{\min}^{\prime}$, since $\widetilde{F}_{\min}$ can be arbitrarily close to 00. For instance, consider the scalar polynomial $F(x)=(x-\epsilon)^{2}+1$ and $K=\mathbb{R}$. Note that for $x_{0}^{2}+x^{2}=1$, we have $\widetilde{F}(x_{0},x)=(x-\epsilon\cdot x_{0})^{2}+x_{0}^{2}=1+\epsilon^{2}\cdot x_{0}^{2}-2\epsilon\cdot x_{0}\cdot x$. Let $x_{0}=\sin\theta$, $x=\cos\theta$. Then, it follows that
$\begin{array}[]{ll}\widetilde{F}(x_{0},x)=1+\epsilon^{2}\cdot\sin^{2}\theta-2\cdot\epsilon\cdot\sin\theta\cos\theta&=1+\frac{\epsilon^{2}}{2}\cdot(1-\cos(2\theta))-\epsilon\cdot\sin(2\theta)\\ &=1+\frac{\epsilon^{2}}{2}-\sqrt{\epsilon^{2}+\frac{\epsilon^{4}}{4}}\sin(2\theta+\phi),\end{array}$
for some $\phi\in[0,2\pi]$. This implies that $\widetilde{F}_{\min}=1+\frac{\epsilon^{2}}{2}-\sqrt{\epsilon^{2}+\frac{\epsilon^{4}}{4}}$. One can verify that $1+\frac{\epsilon^{2}}{2}-\sqrt{\epsilon^{2}+\frac{\epsilon^{4}}{4}}\rightarrow 0$, as $\epsilon\rightarrow\infty$, while $F_{\min}=F_{\min}^{\prime}=1$ for arbitrary $\epsilon$.
The following result is a direct corollary of Theorem 1.2. Its proof is deferred to Section 5.2.
Corollary 1.3.
Let $K$ be as in (1.1) with $m\geq 2$. Suppose that $F(x)$ is an $\ell$$-\ell$ symmetric polynomial matrix with $\deg(F)=d$. If $F\succeq 0$ on $K$, then for any $\varepsilon>0$, we have
provided that $k\geq C\cdot\varepsilon^{-7\eta-3}.$ Here, $\eta$ is the Łojasiewicz exponent in (5.8), and $C>0$ is a constant that depends on $F$, $G$.
Outline of the proofs
@@
Step 1. When $F(x)$ is an $\ell$$-\ell$ symmetric polynomial matrix and $K$ is described by finitely many scalar polynomial inequalities (i.e., $K$ is as in (1.2) ), we establish a polynomial bound on degrees of terms appearing in matrix Putinar’s Positivstellensätz (see Theorem 3.5). It improves the exponential bound given by Helton and Nie [15]. The proof closely follows the idea and strategies of [2, 3], adapting it to the matrix setting. This proof is provided in Section 3 and is completed in three steps:
Step 1a. We perturb $F$ into a polynomial matrix $P$ such that $P$ is positive definite on the scaled simplex $\bar{\Delta}$. To be more specific, we prove that there exist $\lambda>0$ and $h_{i}\in\Sigma[x]$ $(i=1,\dots,m)$ with controlled degrees such that
$P(x):=F(x)-\lambda\sum_{i=1}^{m}h_{i}(x)g_{i}(x)I_{\ell}\succeq\frac{1}{4}F_{\min}\cdot I_{\ell},~{}~{}\forall~{}x\in\bar{\Delta}.$
This step uses a result by Baldi-Mourrain-Parusinski [3, Proposition 3.2] about the approximation of a plateau by SOS polynomials, and some estimates on the matrix Bernstein norm (see Lemma 3.3).
Step 1b. Using the dehomogenized version of matrix Polya’s Positivstellensätz in terms of the Bernstein basis (see Lemma 3.1), we show that there exist matrices $P_{\alpha}^{(k)}\succ 0$ such that
$P=\sum_{\alpha\in\mathbb{N}^{n}_{k}}P_{\alpha}^{(k)}B_{\alpha}^{k}(x),$
and provide an estimate on $k$.
Step 1c. Note that the factors of $B_{\alpha}^{k}(x)$ satisfy $1-x_{1},\dots,1-x_{n},\sqrt{n}-x_{1}-\cdots-x_{n}\in\mbox{QM}[1-\|x\|_{2}^{2}]_{2}$. We know that $P\in\mbox{QM}[1-\|x\|_{2}^{2}]^{\ell}$. Consequently, a polynomial bound on the degrees of terms appearing in matrix Putinar’s Positivstellensätz can be derived from the assumption that $(1-\|x\|_{2}^{2})\cdot I_{\ell}\in\mbox{QM}[G]^{\ell}$.
@@
Step 3. Finally, we prove Theorem 1.1 by applying the main result from Step 1 (i.e., Theorem 3.5) to $F$ and the equivalent scalar polynomial inequalities for $K$ obtained in Step 2. We refer to Section 5 for details.
@@
2 Preliminaries
2.1 Notations
The symbol $\mathbb{N}$ (resp., $\mathbb{R}$) denotes the set of nonnegative integers (resp., real numbers). Let $\mathbb{R}[x]:=\mathbb{R}\left[x_{1},\ldots,x_{n}\right]$ be the ring of polynomials in $x:=\left(x_{1},\ldots,x_{n}\right)$ with real coefficients and let $\mathbb{R}[x]_{d}$ be the ring of polynomials with degrees $\leq d$. For $\alpha=(\alpha_{1},\ldots,\alpha_{n})\in\mathbb{N}^{n}$ and a degree $d$, denote $|\alpha|\coloneqq\alpha_{1}+\cdots+\alpha_{n}$, $\mathbb{N}_{d}^{n}:=\left\{\alpha\in\mathbb{N}^{n}\mid|\alpha|\leq d\right\}.$ For a polynomial $p$, the symbol $\deg(p)$ denotes its total degree. For a polynomial matrix $P=(P_{ij}(x))\in\mathbb{R}[x]^{\ell_{1}\times\ell_{2}}$, its degree is denoted by $\deg(P)$, i.e.,